Session: Feb 2013 (session 1)
Wavelength: 3.6 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (M.Morsiani at Sun Mar 10 15:55:56 2013) Minor failures: TSYS
Something strange on system setup, abnormal signal level after schedule started, power level return at usual level at end of schedule. We cannot understand where was the problem.

Onsala (Michael Lindqvist at Mon Mar 11 07:30:54 2013) Success:
No known problems.

Medicina (Andrea Orlati at Mon Mar 11 15:19:00 2013) Minor failures:
See GK047B feedback, next time we suggest to allocate time for preliminary checks of telescope pointing and proper attenuation level.

Metsahovi (E.Rastorgueva-Foi at Thu Mar 14 09:36:53 2013) Success:
No known problems

Svetloe (Slobojaninova. at Mon Mar 18 11:30:26 2013) Success: